solved problems regarding torsion

Grade:12th Pass

1 Answers

Rahul Kumar
12 years ago

For one propeller drive shaft, compute the torsional shear stress when it is transmitting a torque of 1.76 kNm. The shaft is a hollow tube having an outside diameter of 60 mm and an inside diameter of 40 mm. Find the stress at both the outer and inner surfaces.

Solution:  Given that

torsional shear stress.JPG

Torque transmitted ‘T’ = 1.76 kNm

Outside diameter of tube D0 = 60 mm

Inside diameter of tube Di = 40 mm

Now, using the torsion equation

    T/j = t/r

where J is polar moment of inertia

For hollow shaft

Torsional stress.JPG 

Torsional stress on outer surface is;

t0 = T/J (D0/2) = (1.76 × 1000 × 1000) * (60/2) / p/32 ((60)4 – (40)4)

→   ti = 34.475 N/mm2               

Thus, the stresses on outer and inner surfaces are 51.713 N/mm2 and 34.475 N/mm2 respectively.